OpenStudy (anonymous):

If un+1=5-2u, U2=3 what is u0? How do I sub u_2 in?

OpenStudy (kc_kennylau):

Which field of Mathematics?

OpenStudy (anonymous):

It doesn't say

OpenStudy (kc_kennylau):

What does un, u, U2 and u0 represent?

OpenStudy (kc_kennylau):

did you mean \(u_n\) as in u_n?

OpenStudy (anonymous):

Yeah didn't know how to do that but

OpenStudy (kc_kennylau):

Did you mean \(\Large u_{n+1}=5-2u_n\)?

OpenStudy (anonymous):

Yeah

OpenStudy (kc_kennylau):

Please use brackets next time :) \[u_2=5−2u_1\]\[u_1=5−2u_0\]

OpenStudy (kc_kennylau):

Now plug in \(u_2\) and find \(u_1\), then plug in \(u_1\) to find \(u_0\) :)

OpenStudy (anonymous):

How do I plug u_2 in but

OpenStudy (kc_kennylau):

\[3=5-2u_1,\]solve for \(u_1\).

OpenStudy (perl):

u1 = 1

OpenStudy (perl):

u0 = 2

OpenStudy (anonymous):

Thanks